OpenStudy (anonymous):

How many different non-congruent isosceles triangles can be formed by connecting three of the dots in a 4\times4 square array of dots like the one shown below?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

An array like this?\[\begin{matrix} \cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot \end{matrix}\]

OpenStudy (anonymous):

Let's start by fixing a vertex in (1,1): \[\begin{matrix} \color{red}\bullet&\rightarrow&\color{red}\bullet&&\cdot&&\cdot\\ \uparrow&\swarrow\\ \color{red}\bullet&&\cdot&&\cdot&&\cdot\\ \\ \cdot&&\cdot&&\cdot&&\cdot\\ \\ \cdot&&\cdot&&\cdot&&\cdot \end{matrix}\] \[\begin{matrix} \color{red}\bullet&\rightarrow&\bullet&\rightarrow&\color{red}\bullet&&\cdot\\ \uparrow&&&\swarrow\\ \bullet&&\bullet&&\cdot&&\cdot\\ \uparrow&\swarrow\\ \color{red}\bullet&&\cdot&&\cdot&&\cdot\\ \\ \cdot&&\cdot&&\cdot&&\cdot \end{matrix}\] \[\begin{matrix} \color{red}\bullet&\rightarrow&\bullet&\rightarrow&\bullet&\rightarrow&\color{red}\bullet\\ \uparrow&&&&&\swarrow\\ \bullet&&\cdot&&\bullet&&\cdot\\ \uparrow&&&\swarrow\\ \bullet&&\bullet&&\cdot&&\cdot\\ \uparrow&\swarrow\\ \color{red}\bullet&&\cdot&&\cdot&&\cdot \end{matrix}\] You thus have up to three 45-45-90 triangles for each vertex. For a general counting strategy, you could try fixing two of the red vertices and counting how many vertices you can pick to make up an isosceles triangle.